OpenStudy (anonymous):

calculus help! find an equation of the tangent to the graph of y=arcsin(5x) at the point (1/5(Sq. root of 2), pi/4)

OpenStudy (anonymous):

ick

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

Satellite, please, help here: http://openstudy.com/study#/updates/507cd0dfe4b07c5f7c1fb4c3

OpenStudy (anonymous):

the answer is supposed to be y=5(sq rt 2x)+pi/4-1

OpenStudy (anonymous):

derivative is easy enough, it is \[\frac{5}{\sqrt{1-25x^2}}\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

replace \(x\) by \(\frac{1}{5\sqrt{2}}\) and see what you get for the slope

OpenStudy (anonymous):

can u explain this to be im kinda lost :l

OpenStudy (anonymous):

how to get the derivative, or why to evaluate ?

OpenStudy (anonymous):

how to get the derivative

OpenStudy (anonymous):

:l

OpenStudy (anonymous):

1) memorize the derivative of arcsine it is \(\frac{1}{\sqrt{1-x^2}}\) 2) use the chain rule

OpenStudy (anonymous):

i plug in 5x for x right?

OpenStudy (anonymous):

so since you have \[\arcsin(5x)\] you replace \(x\) by \(5x\) and then multiply by the derivative of \(5x\) which is just 5 that gives you \[\frac{1}{\sqrt{1-(5x)^2}}\times 5=\frac{5}{\sqrt{1-25x^2}}\]

OpenStudy (anonymous):

crystal clear! :}

OpenStudy (anonymous):

actually first expression is just as easy to evaluate as the second so you can leave it like that rest is more or less routine right? find the slope by evaluating the derivative at the first coordinate use the point slope formula finito

OpenStudy (anonymous):

beautiful! thanks so much!

OpenStudy (anonymous):

wait but im not getting what im suppose to get :[